OpenStudy (anonymous):

-x2 + 5 = (4x - 1)2 What is this quadratic equation written in standard form? A. 17x2 - 8x - 4 = 0 B. 17x2 - 8x + 6 = 0 C. 15x2 - 8x + 6 = 0 D. 15x2 - 8x - 4 = 0

OpenStudy (anonymous):

i'm assuming you mean -x^2 +5 = (4x-1)^2 ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

to do this, just expand the right side by doing (4x-1)*(4x-1) which = 16x^2-8x+1 then add that to the left side so (-x^2+5) + (16x^2-8x+1) which = 15x^2-8x+6... so your answer is C

OpenStudy (anonymous):

thanks